For each binary operation βˆ— defined below say whether the following is


a group or not


a) Define βˆ— π‘œπ‘› 𝑍 by π‘Ž βˆ— 𝑏 = π‘Ž βˆ’ 𝑏


b) Define βˆ— π‘œπ‘› 𝑍 by π‘Ž βˆ— 𝑏 = π‘Žπ‘


c) Define βˆ— π‘œπ‘› 𝑅+by π‘Ž βˆ— 𝑏 = π‘Žπ‘


d) Define βˆ— π‘œπ‘› 𝑄 by π‘Ž βˆ— 𝑏 = π‘Žπ‘

1
Expert's answer
2022-03-19T02:39:16-0400
  1. It is not a group, as this law is not associative : "a-(b-c)=a-b+c\neq (a-b)-c" whenever "c\neq 0",
  2. It is not a group, as not every element admits an inverse in "\mathbb{Z}" (for example for "2\in \mathbb{Z}" there is no "n\in \mathbb{Z}" such that "2n=1"),
  3. It is a group, as the group law is associative (the multiplication in "\mathbb{R}" is associative), there is a unity "1\in \mathbb{R}^+" and every element admits an inverse (as there is an inverse in "\mathbb{R}" and for "a>0" the inverse "a^{-1}>0", so it is also in "\mathbb{R}^+" ),
  4. This law is associative and there is a unity, however it is not a group, as the element "0\in\mathbb{Q}" does not admit an element "p\in\mathbb{Q}" such that "0\cdot p =1", so it is not invertible.